# CR-Net: Scaling Parameter-Efficient Training with Cross-Layer Low-Rank Structure

### v1

@@ 1 Introduction
(O1). Low-rank parameter. Representative approaches such as Low-Rank Adaptation (LoRA) and its variants [26, 37, 63, 42, 60, 32] utilize learnable low-rank parameter matrices as memory-efficient substitutes for full-rank weight updates. This design achieves significant parameter reduction while preserving model capacity. Furthermore, this paradigm demonstrates strong compatibility with sparsity-inducing techniques (e.g., SLTrain [19] and LOST [36]) to achieve improved model performance under constrained parameter budgets.
(O2). Low-rank gradient. Recent research has explored the intrinsic low-rank properties of LLM gradients [64, 8, 21, 27, 53, 23, 38]. By projecting optimizer states into low-dimensional subspaces, these methods achieve significant memory reduction during pre-training. Extensions like RSO [11] provide theoretical convergence guarantees under low-rank gradient constraints. Advanced approaches such as Apollo [66] demonstrate additional benefits by combining channel-wise adaptive learning with low-rank gradients, yielding both improved validation accuracy and optimized memory efficiency throughout the pre-training process.
@@
2 Preliminary and Related Works
@@ 2.1 Preliminary
Notations. In this paper, we assume that the amount of transformer blocks of the model is $L$. Denote $s$ and $h$ as the sequence length and hidden dimension, respectively. $h_{\text{ff}}$ stands for the intermediate dimension of the model. We use $X_{l}\in\mathbb{R}^{s\times h}$ to represent the input of layer $l$. $W_{l}^{\text{P}}$, $X_{l}^{\text{P}}$ and $Y_{l}^{\text{P}}$ stand for the parameters, inputs and outputs of linear layers at the position P, where P stands for the detailed position in the transformer layer. For example, $\text{P}\in\{\text{Q},\text{K},\text{V},\text{O},\text{gate},\text{up},\text{down}\}$ in LLaMA-based model and $\text{P}\in\{\text{Q},\text{K},\text{V},\text{O},\text{up},\text{down}\}$ in GPT-based model. Finally, we use $\text{LR}_{r}(A)$ to denote the optimal approximation of $A$ with rank $r$ under Frobenius loss as
@@
$X_{l+1}=\left(\text{silu}(\text{Att}_{l}\cdot W_{l}^{\text{gate}})\odot(\text{Att}_{l}\cdot W_{l}^{\text{up}})\right)W_{l}^{\text{down}}+\text{Att}_{l},$ (2)
@@ 3.1 Observation: previous-Layer and low-Rank term present a better activation approximation
Existing studies have demonstrated that the activations in LLMs exhibit intrinsic low-rank structural properties [12, 29, 62, 42] and thus low-rank approximations of activation matrices achieve acceptable reconstruction fidelity, making pre-training a model with low-rank parameters possible. Furthermore, empirical evidence from [40, 5] highlights inter-layer activation correlations, suggesting that historical activations from preceding layers may encode critical information for current layer computations. Different from those existing observations, we have found a critical structural characteristic that the difference between activations of adjacent layers exhibit a significant low-rank properties.
@@
Emprical evaluation for different activation approximation approached. To compare the performance for different methods including Eq. (3) and direct low-rank approximation for the activation estimation, we conducted comparative experiments by fine-tuning the pre-trained LLaMA-3 8B [18], GPT-2-small [50], and models using the TinyShakespeare dataset. In both schemes we use the same low-rank dimension $r=0.25h$ and we quantify approximation quality via relative error defined as follow.
@@
Figure 2 illustrates the average relative error of two approaches for activation approximation over all the transformer layers. It can be observed that the previous-layer-based yields smaller relative errors, indicating that performing low-rank approximation with the history information provides better reconstruction capability compared to direct approximation of the original activations when preserving the same rank $r$ of singular values. Additional evaluation can be fined in Appendix D.1. We also present a theoretical insight for this observation, whose detail is in Appendix C.
@@ 3.2 Model structure
$Y_{l}^{\text{P}}=\begin{cases}X_{l}^{\text{P}}W_{l}^{\text{P}},\quad l=1,\\ \text{sign}(\beta_{l}^{\text{P}})(|\beta_{l}^{\text{P}}|+\varepsilon)Y_{l-1}^{\text{P}}+X_{l}^{\text{P}}A_{l}^{\text{P}}B_{l}^{\text{P}},\quad l=2,3,\cdots,L.\end{cases}$ (6)
@@ 4 Complexity analysis
In this section, we present the analysis for the parameter complexity, memory complexity, and computation complexity of CR-Net based on LLaMA architecture [58, 18] and Adam optimizer [34]. We also present the communication overhead and HBM memory analysis for CR-Net under multiple GPU training in Appendix F.
@@ 4.1 Complexity analysis without re-computation
Computational Complexity. For the first transformer layer of CR-Net, the forward- and backward- propagation process maintains identical computational complexity to full-rank training. In subsequent layers, CR-Net employs low-rank weights while eliminating the nonlinear activation step applied to low-rank outputs, combined with cross-layer residual connections in contrast to CoLA [42]. Given that both operations introduce lower-order computational complexity terms, we conclude that these layers maintain equivalent FLOPs to CoLA. Following existing analysis [42], the total computation complexity is (see Appendix A.2 for details):
@@ 4.2 Complexity analysis with re-computation
We evaluate recomputation overhead and activation memory costs for CR-Net based on LLaMA-like architecture in B. Table 2 compares activation memory and recomputation complexity across frameworks. including practical measurements for LLaMA2-7B (batch size 16). Althrough our re-computation strategy introduce additional memory overhead than Vanilla GCP and CoLA-M, CR-Net also achieves a 67.4% reduction in total computation overhead compared to Vanilla GCP and a 8.0% reduction compared to CoLA with significant memory saves, validating the effectiveness of our proposed cross-layer framework and re-computation strategy.
Table 2: Activation memory and re-computation complexity of different activation-efficient approaches based on LLaMA architecture with batch size 1 and BF16 precision. For CR-Net, the notation $b$ represents the elements of $\mathcal{A}$ where $b=4$ . The last two column is the evaluated computation and memory of LLaMA2-7B framework with $r=512$, batch size $B=16$, and sequence length $s=256$. We also add a explanation of the re-computation complexity of CoLA-M in Appendix B.2.
@@ 5 Experiments
This section conducts numerical experiments to systematically assess the efficiency of the proposed CR-Net methodology. Our evaluation framework encompasses pre-training tasks spanning diverse model scales. Furthermore, we empirically validate the performance advantages of our innovative re-computation approach. To substantiate design choices, comprehensive ablation studies are performed focusing on two critical components: the determination of optimal rank values for low-rank parameterization and the adaptive learning dynamics of scaling coefficients $\beta_{l}^{\text{P}}$ within our parameterization scheme. More experimental results can be referred in Appendix D.
Table 3: Comparsion of validation perplexity (PPL) ($\downarrow$), parameter complexity (M) ($\downarrow$) and evaluated memory overhead (GB) ($\downarrow$) of parameter, gradients, and optimizer states of different effective training approaches in the pre-training task of LLaMA model with C4-en dataset. The results of compared methods are referred from [27, 19, 42, 11, 66]. For CR-Net, it is compared with other parameter-efficient methods with aligned parameter complexity (marked as $\Diamond$) and compared with other optimizer-effiencnt methods with aligned memory overhead (marked as $\dagger$).
@@
Table 4: Comparsion of validation perplexity ($\downarrow$) and memory ($\downarrow$) of different approaches in LLaMA-2 7B pre-training tasks. The results of compared methods are referred from [42, 66].
Memory (GB) 10K 40K 65K 80K
@@ 5.1 Pretraining with CR-Net
Experiment setup. We evaluate the proposed CR-Net framework by pre-training LLaMA-2 models [58] with size varying from 60M to 7B. Following the existing experimental settings [64], we train the model over C4-en dataset [52], which is primarily intended for pre-training language models and word representations with large scale. We compared CR-Net with existing parameter-efficient methods including LoRA [26], ReLoRA [37], SLTrain [19], CoLA [42] with aligned parameter complexity. We also compared CR-Net with other optimizer-efficient algorithms including GaLore [64], RSO [11], and Apollo [66] with aligned memory overhead. We use the same configurations as those reported in [64], and the detailed experiment setting are in Appendix D.2.1.
@@
Figure 6 illustrates the through of ours when pretraining LLaMA-2 1B model on a A100 40G GPU with batch size 16, from which one can observe that CR-Net achieve the best throughout compared with other methods. Even when considering the data-parallel communication overhead across 4 GPUs, CR-Net achieves over 6% improvement compared to state-of-the-art methods without such communication.
@@
Figure 3: The throughout of algorithms for training LLaMA-2 1B on a Nvidia A100 40G GPU. The result of other comparable methods are from [42].
@@
Figure 5: The comparsion of evaluation perplexity for CR-Net in training LLaMA-2 350M with fixed $\beta_{l}^{\text{P}}$ and learnable $\beta_{l}^{\text{P}}$.
@@
5.2 Ablations
How does rank selection impact pre-training performance? We train the LLaMA-2 350B network with CR-Net using identical parameters but varying ranks across transformer layers. Experimental details are provided in Appendix D.2.5. Figure 6 shows that networks with higher ranks in middle layers and lower ranks in side layers exhibit superior performance.
Whether does the learnable scaling factor $\beta_{l}^{\text{P}}$ benefit the model convergence? We train the LLaMA-2 350M network with the same hyperparameters as that in Section 5.1 except make the scaling $\beta_{l}^{\text{P}}$. From the perplexity shown in Figure 6, we can obtain that training CR-Net with a learnable $\beta_{l}^{\text{P}}$ can improve the numerical stability as there are spikes in the runs with $\beta_{l}^{\text{P}}$ fixed. This match our ovservation in Section that the best $\beta_{l}^{\text{P}}$ to achieve the lowest rank varies during the training process. More discussions can be found in Appendix E.2.
Whether other cross-layer residual strategies do well in pre-training with low-rank parameters? We train the LLaMA-2 1B network with low-rank parameter for 50000 iterations in all linear layers except the first transformer layer with CR-Net , Learnable-ResFormer [65] and DenseFormer [49]. Figure 6 illustrates the evaluation perplexity of both cross-layer residual architectures. It can be observation that CR-Net outperforms the ResFormer architecture in pre-training with low-rank parameters.
@@ 6 Conclusion
Appendix A Computation and memory analysis of CR-Net without re-computation
@@
A.1 Transformer layers with LLaMA architecture
In order to make a comparsion with other efficient methods, the computation and memory analysis are based on LLaMA framework [58] with $L$ transformer layers. Here, we focus on the forward- and backward-propagation of CR-Net. It should be reminded that we omit the layer normalization, rotary position embedding, and in-layer residual for simply.
@@ A.1 Transformer layers with LLaMA architecture
$Y_{1}^{\text{gate}}=Y_{1}^{\text{O}}W_{1}^{\text{gate}},\quad Y_{1}^{\text{up}}=Y_{1}^{\text{O}}W_{1}^{\text{up}},\quad X_{1}^{\text{down}}=\text{silu}(Y_{1}^{\text{gate}})\odot Y_{1}^{\text{up}},$ (16)
@@
$Y_{l}^{\text{gate}}=\beta_{l}^{\text{gate}}Y_{l-1}^{\text{gate}}+Y_{l}^{\text{O}}A_{l}^{\text{gate}}B_{l}^{\text{gate}},\quad Y_{l}^{\text{up}}=\beta_{l}^{\text{up}}Y_{l-1}^{\text{up}}+Y_{l}^{\text{O}}A_{l}^{\text{up}}B_{l}^{\text{up}},\quad X_{l}^{\text{down}}=\text{silu}(Y_{l}^{\text{gate}})\odot Y_{l}^{\text{up}},$ (17)
@@
$\mathcal{D}Y_{L}^{\text{up}}=\mathcal{D}X_{L}^{\text{down}}\odot\text{silu}(Y_{L}^{\text{gate}}),\quad\mathcal{D}Y_{L}^{\text{gate}}=\mathcal{D}X_{L}^{\text{down}}\odot Y_{L}^{\text{up}}\odot\text{silu}^{\prime}(Y_{L}^{\text{gate}}).$ (23)
@@
$\mathcal{D}Y_{l}^{\text{up}}$ $=\beta_{l+1}^{\text{up}}\mathcal{D}Y_{l+1}^{\text{up}}+\mathcal{D}X_{l}^{\text{down}}\odot\text{silu}(Y_{l}^{\text{gate}}),$ (24)
$\mathcal{D}Y_{l}^{\text{gate}}$ $=\beta_{l+1}^{\text{gate}}\mathcal{D}Y_{l+1}^{\text{gate}}+\mathcal{D}X_{l}^{\text{down}}\odot Y_{l}^{\text{up}}\odot\text{silu}^{\prime}(Y_{l}^{\text{gate}}).$
@@
A.2 Computation analysis without re-computation

### v2

@@ 1 Introduction
(O1). Low-rank parameter. Representative approaches such as Low-Rank Adaptation (LoRA) and its variants [27, 38, 66, 43, 63, 33, 47, 15] utilize learnable low-rank parameter matrices as memory-efficient substitutes for full-rank weight updates. This design achieves significant parameter reduction while preserving model capacity and can achieve further memory savings with quantized parameters [15]. Furthermore, this paradigm demonstrates strong compatibility with sparsity-inducing techniques (e.g., SLTrain [20] and LOST [37]) to achieve improved model performance under constrained parameter budgets.
(O2). Low-rank gradient. Recent research has explored the intrinsic low-rank properties of LLM gradients [67, 8, 22, 28, 56, 24, 39]. By projecting optimizer states into low-dimensional subspaces, these methods achieve significant memory reduction during pre-training. Extensions like RSO [11] provide theoretical convergence guarantees under low-rank gradient constraints. Advanced approaches such as Apollo [69] demonstrate additional benefits by combining channel-wise adaptive learning with low-rank gradients, yielding both improved validation accuracy and optimized memory efficiency throughput the pre-training process.
@@
2 Related Works and Preliminary
2.1 Related works
Here we display some prior works of parameter-efficient training for LLMs. More related works can be found in Appendix A.
@@ 2.2 Preliminary
Notations. In this paper, we assume that the amount of transformer blocks of the model is $L$. Denote $s$ and $h$ as the sequence length and hidden dimension, respectively. $h_{\text{ff}}$ stands for the intermediate dimension of the model. We use $X_{l}\in\mathbb{R}^{s\times h}$ to represent the input of layer $l$. $W_{l}^{\text{P}}$, $X_{l}^{\text{P}}$ and $Y_{l}^{\text{P}}$ stand for the parameters, inputs and outputs of linear layers at the position P. For example, $\text{P}\in\{\text{Q},\text{K},\text{V},\text{O},\text{gate},\text{up},\text{down}\}$ in LLaMA-based model and $\text{P}\in\{\text{Q},\text{K},\text{V},\text{O},\text{up},\text{down}\}$ in GPT-based model. Finally, we use $\text{LR}_{r}(A)$ to denote the optimal approximation of $A$ with rank $r$ under Frobenius loss as
@@
$X_{l+1}=\left(\text{{SwiGLU}}(\text{Att}_{l}\cdot W_{l}^{\text{gate}})\odot(\text{Att}_{l}\cdot W_{l}^{\text{up}})\right)W_{l}^{\text{down}}+\text{Att}_{l},$ (2)
@@ 3.1 Observation: previous-Layer and low-Rank term present a better activation approximation
Existing studies have demonstrated that the activations in LLMs exhibit intrinsic low-rank structural properties [12, 30, 65, 43] and thus low-rank approximations of activation matrices achieve acceptable reconstruction fidelity, making pre-training a model with low-rank parameters possible. Furthermore, empirical evidence from [41, 5] highlights inter-layer activation correlations, suggesting that historical activations from preceding layers may encode critical information for current layer computations. Unlike existing observations, we have found a critical structural characteristic that the difference between activations of adjacent layers exhibits a significant low-rank property.
@@
Emprical evaluation for different activation approximation approaches. To compare the performance for different methods including Eq. (3) and direct low-rank approximation for the activation estimation, we conducted comparative experiments by fine-tuning the pre-trained LLaMA-3 8B [19], GPT-2-small [53], and models using the TinyShakespeare dataset. In both schemes we use the same low-rank dimension $r=0.25h$ and we quantify approximation quality via relative error defined as follow.
@@
Figure 2 illustrates the average relative error of two approaches for activation approximation over all the transformer layers. It can be observed that the previous-layer-based yields smaller relative errors, indicating that performing low-rank approximation with the history information provides better reconstruction capability compared to direct approximation of the original activations when preserving the same rank $r$ of singular values. Additional evaluation can be fined in Appendix E.1. We also present a theoretical insight for this observation, whose detail is in Appendix D.
@@ 3.2 Model structure
$Y_{l}^{\text{P}}=\begin{cases}X_{l}^{\text{P}}W_{l}^{\text{P}},\hskip 126.61467ptl=1,\\ \text{sign}(\beta_{l}^{\text{P}})(|\beta_{l}^{\text{P}}|+\varepsilon)Y_{l-1}^{\text{P}}+X_{l}^{\text{P}}A_{l}^{\text{P}}B_{l}^{\text{P}},\quad l=2,3,\cdots,L.\end{cases}$ (6)
@@
CR-Net’s Stability Mechanism. CR-Net overcomes the instability of low-rank pretraining by preserving high-rank activations through a full-rank first layer and modulated residual connections. A learnable scalar $\beta$ dynamically balances the high-rank state and low-rank increments, enabling robust signal reconstruction without collapsing into low-dimensional subspaces. Unlike methods that enforce strict low-rank constraints via projections or decompositions (e.g., QR, SVD), which often cause information loss and numerical issues, CR-Net uses standard optimization methods without additional overhead. This design ensures stable training dynamics, similar to full-rank approaches, while achieving significant parameter and memory reductions.
@@ 4 Complexity analysis
In this section, we present the analysis for the parameter complexity, memory complexity, and computation complexity of CR-Net based on LLaMA architecture [61, 19] and Adam optimizer [35]. We also present the communication overhead and HBM memory analysis for CR-Net under multiple GPU training in Appendix G.
@@ 4.1 Complexity analysis without re-computation
Computational Complexity. For the first transformer layer of CR-Net, the forward- and backward- propagation process maintains identical computational complexity to full-rank training. In subsequent layers, CR-Net employs low-rank weights while eliminating the nonlinear activation step applied to low-rank outputs, combined with cross-layer residual connections in contrast to CoLA [43]. Given that both operations introduce lower-order computational complexity terms, we conclude that these layers maintain equivalent FLOPs to CoLA. Following existing analysis [43], the total computation complexity is (see Appendix B.2 for details):
@@ 4.2 Complexity analysis with re-computation
We evaluate recomputation overhead and activation memory costs for CR-Net based on LLaMA-like architecture in C. Table 2 compares activation memory and recomputation complexity across frameworks. including practical measurements for LLaMA2-7B (batch size 16). Althrough our re-computation strategy introduce additional memory overhead than Vanilla GCP and CoLA-M, CR-Net also achieves a 67.4% reduction in total computation overhead compared to Vanilla GCP and a 8.0% reduction compared to CoLA with significant memory saves, validating the effectiveness of our proposed cross-layer framework and re-computation strategy.
Table 2: Activation memory and re-computation complexity of different activation-efficient approaches based on LLaMA architecture with batch size 1 and BF16 precision. For CR-Net, the notation $b$ represents the elements of $\mathcal{A}$ where $b=4$ . The last two column is the evaluated computation and memory of LLaMA2-7B framework with $r=512$, batch size $B=16$, and sequence length $s=256$. We also add a explanation of the re-computation complexity of CoLA-M in Appendix C.2.
@@ 5 Experiments
This section conducts numerical experiments to systematically assess the efficiency of the proposed CR-Net methodology. Our evaluation framework encompasses pre-training tasks spanning diverse model scales. Furthermore, we empirically validate the performance advantages of our innovative re-computation approach. To substantiate design choices, comprehensive ablation studies are performed focusing on two critical components: the determination of optimal rank values for low-rank parameterization and the adaptive learning dynamics of scaling coefficients $\beta_{l}^{\text{P}}$ within our parameterization scheme. More experimental results can be referred in Appendix E.
Table 3: Comparison of validation perplexity (PPL) ($\downarrow$), parameter complexity (M) ($\downarrow$) and evaluated memory overhead (GB) ($\downarrow$) of parameter, gradients, and optimizer states of different effective training approaches in the pre-training task of LLaMA model with C4-en dataset. The results of compared methods are referred from [28, 20, 43, 11, 69, 47, 48]. For CR-Net, it is compared with other parameter-efficient methods with aligned parameter complexity (marked as $\Diamond$) and compared with other optimizer-effiencnt methods with aligned memory overhead (marked as $\dagger$). N.A. means that the corresponding experiment has not been taken.
@@
Table 4: Comparison of validation perplexity ($\downarrow$) and memory ($\downarrow$) of different approaches in LLaMA-2 7B pre-training tasks. The results of compared methods are referred from [43, 69].
Memory (GB) Training steps
10K 40K 65K 80K
@@
Table 5: Comparison of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 13B pre-training tasks.
Steps 40K
8-bit Adam 17.85
CR-Net w. re-computation 18.12
Table 6: Comparison of validation perplexity ($\downarrow$) for CR-Net with or without low-rank first layer in LLaMA-2 350M pre-training tasks.
Training tokens 6.4B
CR-Net w.o. low-rank first layer 18.95
CR-Net w. low-rank first layer 19.68
@@ 5.1 Pretraining with CR-Net
Experiment setup. We evaluate the proposed CR-Net framework by pre-training LLaMA-2 models [61] with size varying from 60M to 13B. Following the existing experimental settings [67], we train the model over C4-en dataset [55], which is primarily intended for pre-training language models and word representations with large scale. We compared CR-Net with existing parameter-efficient methods including LoRA [27], ReLoRA [38], VeLoRA [47], FLoRA [22], SLTrain [20], SLTrain [20], CoLA [43], LORO [48] with aligned parameter complexity. We also compared CR-Net with other optimizer-efficient algorithms including GaLore [67], RSO [11], and Apollo [69] with aligned memory overhead. We use the same configurations as those reported in [67], and the detailed experiment setting are in Appendix E.2.1. For models with sizes ranging from 60M to 1B parameters, we do not employ the re-computation strategy to maintain consistency with existing baselines. For the 7B and 13B models, we employ re-computation to reduce memory consumption.
@@
Furthermore, we report the validation perplexity for the pre-training of LLaMA-2 13B after 40,000 initialization steps. Experimental details are provided in Appendix E.2.3. As demonstrated in Table 6, CR-Net achieves more than 50% reduction in parameters, while incurring only a 2% degradation in validation performance.
Pre-training and inference throughput. Figure 3 illustrates the throughput performance of our method alongside comparable approaches in both pre-training and inference tasks, with detailed experimental configurations provided in Appendix E.2.1. As shown, CR-Net achieves superior throughput compared to other methods in LLaMA-2 1B pre-training and inference scenarios. Notably, even when accounting for data-parallel communication overhead across 4 GPUs for LLaMA-2 1B pre-training, CR-Net demonstrates over 6% improvement relative to state-of-the-art methods that do not incur such communication costs. Furthermore, CR-Net exhibits an 8.6% throughput enhancement compared to LORO in LLaMA-2 7B pre-training.
5.2 Ablations
How does rank selection impact pre-training performance? We train the LLaMA-2 350B network with CR-Net using identical parameters but varying ranks across transformer layers. Experimental details are provided in Appendix E.2.6. Figure 6 shows that networks with higher ranks in middle layers and lower ranks in side layers exhibit superior performance.
Whether does the learnable scaling factor $\beta_{l}^{\text{P}}$ benefit the model convergence? We train the LLaMA-2 350M network with the same hyperparameters as that in Section 5.1 except make the scaling $\beta_{l}^{\text{P}}$. From the perplexity shown in Figure 6, we can obtain that training CR-Net with a learnable $\beta_{l}^{\text{P}}$ can improve the numerical stability as there are spikes in the runs with $\beta_{l}^{\text{P}}$ fixed. This match our observation in Section that the best $\beta_{l}^{\text{P}}$ to achieve the lowest rank varies during the training process. More discussions can be found in Appendix F.2.
Whether other cross-layer residual strategies do well in pre-training with low-rank parameters? To compare different cross-layer residual patterns under the same global low-rank parameterization, we implement a Learnable‑ResFormer variant that replaces CR-Net’s previous-layer residual term with a learnable skip from the first layer’s activation across all low-rank linear layers, inspired by the first-layer value residual in ResFormer [68] and DenseFormer [52]. As Figure 6 illustrates, CR-Net outperforms the other two low-rank adaptation under the similar low-rank parameter budget.
Whether the full-rank first layer is necessary? We trained the LLaMA-2 350M model with either full-rank or low-rank parameters in the first transformer layer. As shown in Table 6, utilizing a low-rank first layer introduces a 3.5% degradation in validation perplexity. This finding demonstrates the critical role of the full-rank first layer in reconstructing high-rank signals from low-rank information.
Figure 3: The average throughput (tokens/s) for each device of different algorithms. (Left: LLaMA-2 1B pre-training on an Nvidia A100 40G GPU, with results of other comparable methods from [43]. Middle: LLaMA-2 1B inference on a Nvidia A100 80G GPU. Right: LLaMA-2 7B pre-training on a Nvidia A100 80G GPU.)
@@ 5.2 Ablations
Figure 5: The Comparison of evaluation perplexity for CR-Net in training LLaMA-2 350M with fixed $\beta_{l}^{\text{P}}$ and learnable $\beta_{l}^{\text{P}}$.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Key Research and Development Program of China (No. 2024YFA1012902) and National Natural Science Foundation of China (No. 12288101, 92370121, 12301392). This work is also supported by AI for Science Institute, Beijing, China.
@@ Acknowledgments
Appendix A Additional related works
Appendix B Computation and memory analysis of CR-Net without re-computation
@@
B.1 Transformer layers with LLaMA architecture
In order to make a Comparison with other efficient methods, the computation and memory analysis are based on LLaMA framework [61] with $L$ transformer layers. Here, we focus on the forward- and backward-propagation of CR-Net. It should be reminded that we omit the layer normalization, rotary position embedding, and in-layer residual for simply.
@@ B.1 Transformer layers with LLaMA architecture
$Y_{1}^{\text{gate}}=Y_{1}^{\text{O}}W_{1}^{\text{gate}},\quad Y_{1}^{\text{up}}=Y_{1}^{\text{O}}W_{1}^{\text{up}},\quad X_{1}^{\text{down}}=\text{{SwiGLU}}(Y_{1}^{\text{gate}})\odot Y_{1}^{\text{up}},$ (16)
@@
$Y_{l}^{\text{gate}}$ $=\beta_{l}^{\text{gate}}Y_{l-1}^{\text{gate}}+Y_{l}^{\text{O}}A_{l}^{\text{gate}}B_{l}^{\text{gate}},\quad Y_{l}^{\text{up}}=\beta_{l}^{\text{up}}Y_{l-1}^{\text{up}}+Y_{l}^{\text{O}}A_{l}^{\text{up}}B_{l}^{\text{up}},$ (17)
$X_{l}^{\text{down}}$ $=\text{{SwiGLU}}(Y_{l}^{\text{gate}})\odot Y_{l}^{\text{up}},$
@@
$\mathcal{D}Y_{L}^{\text{up}}=\mathcal{D}X_{L}^{\text{down}}\odot\text{{SwiGLU}}(Y_{L}^{\text{gate}}),\quad\mathcal{D}Y_{L}^{\text{gate}}=\mathcal{D}X_{L}^{\text{down}}\odot Y_{L}^{\text{up}}\odot\text{{SwiGLU}}^{\prime}(Y_{L}^{\text{gate}}).$ (23)
@@
$\mathcal{D}Y_{l}^{\text{up}}$ $=\beta_{l+1}^{\text{up}}\mathcal{D}Y_{l+1}^{\text{up}}+\mathcal{D}X_{l}^{\text{down}}\odot\text{{SwiGLU}}(Y_{l}^{\text{gate}}),$ (24)
$\mathcal{D}Y_{l}^{\text{gate}}$ $=\beta_{l+1}^{\text{gate}}\mathcal{D}Y_{l+1}^{\text{gate}}+\mathcal{D}X_{l}^{\text{down}}\odot Y_{l}^{\text{up}}\odot\text{{SwiGLU}}^{\prime}(Y_{l}^{\text{gate}}).$
@@
B.2 Computation analysis without re-computation